OpenStudy (superhelp101):

How many grams of Ca are needed to react completely with 2.20 L of a 4.50 M HCl solution? Ca + 2HCl yields CaCl2 + H2 5.00 g 9.90 g 198 g 361 g

OpenStudy (superhelp101):

@Abmon98

OpenStudy (abmon98):

Number of Moles=Concentration*Volume Number of Moles of Ca(s)=2.20*4.50=9.9 moles Mole Ratio of Ca(s) to HCl 1 mole of Ca:2 mole of HCl x mole of Ca :9.9 mole of HCl x=9.9*1/2=4.95 moles of Ca Number of moles=Mass(g)/Molar Mass(g/mol) Mass(g)=Molar Mass(g/mol)*Number of Moles Mass(g)=40.1*4.95=198.495 grams

OpenStudy (superhelp101):

WOW тн@ηк ⑂øü‼

OpenStudy (abmon98):

Your most welcome :D :)

OpenStudy (somy):

u did the whole working again <.< @Abmon98

OpenStudy (abmon98):

@somy Lol i forgot :P

OpenStudy (somy):

@Abmon98 i'll forget you next time i catch you >_< u r smarty pants so use your knowledge wisely - helping other's actually learn :3 believe me - if you see someone you taught couple of days or whatever ago, doing the question on their own or helping someone else with the thing U taught u will feel happy :)

OpenStudy (abmon98):

@somy hahaha you are the smarty one i thought by writing out steps and the answer directly is a way of teaching others a method to solve their question and could help them in learning but looks like its not :/